OpenStudy (anonymous):

Solve by completing the square and then use the square root method to solve. x2 - 4x - 32 = 0

OpenStudy (radar):

A=1, B=-4, and C=-32. To complete the square method, First A must =1, and it does, so we can proceed. Divide B by 2, so we have a value of -2 and square this giving us a 4. Now add this 4 in the following manner: \[(x ^{2}-4x + 4)-32 -4\] Note we added 4 but took it away so we have not changed the value of the equation We now have a perfect square \[(x ^{2}-4x+4)=(x-2)^{2}\] Now we can write \[(x-2)^{2}=36\]after adding 36 to both sides to keep it legal. Now take the sq rt of both sides getting x-2=+/-6 x=8 x=-4

OpenStudy (radar):

I guess we did the "complete the square" thing followed with the "square root method." But I haven't heard of the "square root method" just completing the square, factoring, and use of the quadratic equation.